ncert class 11 physics  chapter 2 example 2.2 
How to find value of tan 40?

Grade:11

1 Answers

Abdul Rahman
3 years ago
the best way is to use the Maclaurin series

tanx = x + 1x​3/3 + 2x5/15 + 17x7/317 + 62x9/2835.....(x is angle in radians)

x = 40 degrees = 2/9pi = 0.7 radians

tan 40 = 0.7 + (0.7)3/3 + 2(0.7)5/15 + 17(0.7)7/317 +....
tan 40 = 0.7 + 0.11433 + 0.0.0224 + 0.004
tan 40 = 0.84 (approx.)